# Thin gap approximations for microfluidic device design

### v1

@@ Abstract
Over 125 years ago, Henry Selby Hele-Shaw realized that the depth-averaged flow in thin gap geometries can be closely approximated by two-dimensional (2D) potential flow, in a surprising marriage between the theories of viscous-dominated and inviscid flows. Hele-Shaw flows allow visualization of potential flows over 2D airfoils and also undergird important discoveries in the dynamics of interfacial instabilities and convection, yet they have found little use in modeling flows in microfluidic devices, although these devices often have thin gap geometries. Here, we derive a Hele-Shaw approximation for the flow in the kinds of thin gap geometries created within microfluidic devices. Although these equations have been reported before, prior work used a less direct derivation. Here, we obtain them via a modified Method of Weighted Residuals (MWR), interpreting the Hele–Shaw approximation as the leading term of an orthogonal polynomial expansion that can be systematically extended to higher-order corrections. We provide substantial numerical evidence showing that approximate equations can successfully model real microfluidic and inertial-microfluidic device geometries. By reducing three-dimensional (3D) flows to 2D models, our validated model will allow for accelerated device modeling and design.
@@ 1 Introduction
Although Hele-Shaw cells are often thought of as devices for engineering reduced complexity (2D) flows, thin gap domains are also commonly realized in microfluidic devices Stone et al. (2004); Di Carlo (2009), due to the nature of the fabrication methods used to create them. Because the Hele-Shaw approximation can not model no-slip boundary conditions on in-plane boundaries, it has found little use for dissecting the function of existing devices or designing new ones. Yet, it remains unanswered whether a different 2D model from Hele-Shaw exists, allowing the key dynamics present within a microfluidic or inertial microfluidic device to be predicted or understood, at far less computational cost than a fully 3D flow model.
The classical derivation of Hele-Shaw’s approximation assumes that, within the thin-gap geometry, the flow is planar but varies with $z$, the gap coordinate. Specifically assuming gap boundaries at $z=\pm\frac{h}{2}$, our putative velocity field is parabolic in $z$: $\mathbf{u}(x,y,z)\approx(u_{1}(x,y),v_{1}(x,y),0)\,6\,(\frac{1}{4}-\frac{z^{2}}{h^{2}})$, with $u_{1}$, $v_{1}$ the depth averaged velocities. Approximating: $\mu\nabla^{2}\mathbf{u}\approx\mu\partial^{2}\mathbf{u}/\partial z^{2}=-\frac{12\mu}{h^{2}}(u_{1}(x,y),v_{1}(x,y),0)$, and balancing with the pressure gradient, we arrive at a form of Darcy’s equation: $(u_{1},v_{1})=-\frac{h^{2}}{12\mu}\nabla p$, which models the pressure-driven flow of fluid through a porous medium, when velocities are averaged over multiple pores Whitaker (1986). Assuming that the flow is incompressible, $\nabla\cdot\mathbf{u}=0$, we obtain: $\nabla^{2}p=0$, so that the thin gap flow, $\mathbf{u}$, is a scalar potential flow, and $\frac{h^{2}p}{12\mu}$ the harmonic potential. Thus, despite the fact that flow between the plates is viscously dominated, the mean flows are captured by the same potential theory that models high-speed inviscid flows, such as those of ideal fluids over airfoils (Fig. 1(a,b)).
@@
Previous work has aimed to incorporate these additional effects into Hele-Shaw-type models. One approach involves adding boundary layer corrections Balsa (1998), though these calculations are often complex and highly geometry-dependent. Another direction is to modify the governing equations themselves. The Darcy-Brinkman equations add in-plane viscous stresses to Darcy’s law, Brinkman (1949), and can be used for porous media Tsay & Weinbaum (1991); Howells (1974), as well as to thin-gap geometries with empirically determined relative coefficients on the in-plane and in-depth stresses Pepper et al. (2010). There are, by contrast, multiple pathways for incorporating inertial effects. When modeling porous media flow, a nonlinear pressure-flow relation (Darcy-Forchheimer) is commonly imposed Chen et al. (2001); Guta & Sundar (2010). Conversely, when Darcy’s law is arrived at by integrating the flow equations over some thin dimension, inertia-containing terms may likewise be integrated, though there are disagreements between papers on the precise coefficients thereby obtained Desvillettes et al. (2008); Inamdar et al. (2020).
@@
Here, we arrive at a Hele Shaw-like approximation via a systematic expansion of the flow-field in the gap coordinate. Our approximation naturally includes inertial and in-plane viscous stresses. Comparisons with full 3D simulations shows that our approximation accurately models flows in geometries used in real microfluidic devices, including both conventional ($\mathrm{Re}=0$) and inertial ($\mathrm{Re}=1-100$) microfluidic flows.
2 A new perspective of the Hele-Shaw approximation
@@
We consider the Navier–Stokes equations for an incompressible, homogeneous fluid confined within a uniform-thickness gap. The equations are non-dimensionalized using the gap height $h$, so that the vertical coordinate $z$ lies in the interval 1/2<z<1/2-1/2<z<1/2. We expand the velocity field as a series, $\mathbf{u}=\sum_{i=1}^{\infty}\mathbf{u}_{i}(x,y)Q_{i}(z)$. We impose a no-slip boundary condition on $\mathbf{u}_{i}$ on any in-plane boundaries. Because the Hele-Shaw approximation yields a parabolic profile for the in-plane velocity components in the narrow gap limit, we select $Q_{1}(z)=z^{2}-1/4$ as the leading vertical basis function. Moreover, because all velocity components must vanish at the boundaries $z=\pm 1/2$, a natural choice for the basis that captures the vertical variation is the Gegenbauer (ultraspherical) polynomials with parameter $\alpha=-1/2$ Belinsky (2000), shifted to the domain 1/2<z<1/2-1/2<z<1/2: $Q_{n}(z)=\frac{1}{2}\int_{-1}^{2z}P_{n}(\zeta)\mathrm{d}\zeta$, where $P_{n}$ is the $n$-th Legendre polynomial. We obtain a sequence of basis functions: $Q_{1}=(z^{2}-1/4)$, $Q_{2}=2z(z^{2}-1/4)$, $Q_{3}=(5z^{2}-1/4)(z^{2}-1/4)$. Projection of our equations onto this basis benefits from the fact that $\left\langle Q_{m},Q_{n}\right\rangle=0$ for $m\neq 0$, where $\left\langle f,g\right\rangle\equiv\int_{-1/2}^{1/2}f(z)g(z)(1/4-z^{2})^{-1}\mathrm{d}z$. However, when the series expansion of $\mathbf{u}$ is substituted into the Navier-Stokes equations, there is no way to isolate the terms in $Q_{1}$ from higher order polynomial contributions. As a first step toward resolving this closure problem, we consider applying the series expansion to a unidirectional shear flow (rectilinear flow).
@@
Poiseuille flow: We first consider pressure-driven flow in a channel, aligned with the $x$-axis, with uniform rectangular cross-section $(y,z)\in[-L/2,L/2]\times[-1/2,1/2]$. Supposing an imposed pressure gradient of $G$, and scaling $p_{x}$ by $G$, and velocities by $Gh^{2}/\mu$, the steady-state Navier-Stokes equations is $(u(y,z),0,0)$, where $u$ satisfies a 2D Poisson equation $(\partial_{y}^{2}+\partial_{z}^{2})u=-1$. In the rectangular channel, this equation admits a Fourier series solution, first derived by Boussinesq (1868). By substituting the leading order forms of the components of $\mathbf{u}$ into the Poisson equation, we obtain:
@@
where the first two terms come from $\Delta$, and the third term is the dimensionless pressure gradient. Including additional bases such as $Q_{3},Q_{4},\ldots$ in our expansion would allow us to eliminate the $z$-dependent terms in Eq. (1), but at the expense of introducing a coupled system of equations for the unknown functions $u_{i}$. Before considering the more complex coupled system, we first seek an optimal closure involving only $u_{1}$. We assume the optimal closure is of the form: $au_{1}^{\prime\prime}-12u_{1}+1=0$ for some constant, $a$, to be determined. Substituting into Eq. (1) leads to $r(z;a)u_{1}^{\prime\prime}(y)\approx 0$, where $r(z;a)=6\left(\frac{1}{4}-z^{2}\right)-a$. Since $u_{1}^{\prime\prime}(y)=0$ produces only trivial solutions, we choose $a$ to minimize the left-hand side, integrating the squared residual against a weight function $w(z)$:
$E(a;w)\equiv\int\limits_{-\frac{1}{2}}^{\frac{1}{2}}\left(6\left(\frac{1}{4}-z^{2}\right)-a\right)^{2}w(z)\mathrm{d}z.$ (2)
@@
$u_{1}(y)=-\frac{\left(1-\cosh(\sqrt{12}y)\operatorname{sech}\left(L\sqrt{3}\right)\right)}{2}\left(\frac{1}{4}-z^{2}\right).$ (3)
@@
The key step to producing a better approximation is to observe that the space of functions spanned by Gegenbauer polynomials remains the natural space in which to optimize our approximation of Eq. (1). Since terms in Eq. (1) do not vanish at $z=\pm 1/2$, we multiply the equation by a weight $w(z)=(1/4-z^{2})$. Then $r(z;a)w(z)=\sum_{m}\beta_{m}Q_{m}$, for some set of coefficients $\beta_{m}$, and applying Parseval’s theorem to this series expansion gives:
$E(a;w)\equiv\int_{-1/2}^{1/2}r(z;a)^{2}(1/4-z^{2})\mathrm{d}z=\left\langle w(z)r,w(z)r\right\rangle=\sum_{m}\beta_{m}^{2}\left\langle Q_{m},Q_{m}\right\rangle^{2}.$ (4)
@@
Therefore, our optimal closure consists of minimizing Eq. (2) with $w(z)=(1/4-z^{2})$. Under this weight, we obtain $E(a;w)=a^{2}/6-2a/5+9/35$, which is minimized when $a=6/5$. This optimal closure leads to a different optimal profile:
@@
$u_{1}(y)=-\frac{\left(1-\cosh(\sqrt{10}y)\operatorname{sech}\left(L\sqrt{\frac{5}{2}}\right)\right)}{2}\left(\frac{1}{4}-z^{2}\right).$ (5)
@@
Figure 2: Validating a 2D approximation for the pressure-driven flow in a rectangular channel. (a) Flow profile at $z=0$: exact solution (black curve), approximation from Eq. (3) (red curve), and approximation from Eq. (5) (blue dashed curve), for $L=1$. (b) Relative error, measured by the $L^{\infty}$ norm $\frac{\lVert u-u_{\text{app}}\rVert_{\infty}}{\lVert u\rVert_{\infty}}$, as a function of aspect ratio $L$. Red curve: Eq. (3), blue dashed curve: Eq. (5).
Both approximations derived above are compared against the exact solution. The choice $a=6/5$ consistently yields a more accurate approximation across the entire $y$-domain (Fig. 2a) and for all tested values of $L$ (Fig. 2b). For example, when $L=1$, the relative errors for $a=1$ and $a=6/5$ are 0.115 and 0.031, respectively. The accuracy of the $a=6/5$ approximation is particularly noteworthy, given that $L=1$ does not satisfy the classical scale-separation assumption $L\gg 1$. As $L$ increases beyond 1, the relative error decreases rapidly for both approximations, indicating convergence toward the asymptotic limit. To probe the optimality of our closure relation more closely, at each aspect ratio, we numerically compute the error-minimizing value of $a$ (determined by minimizing the relative $L^{\infty}$ error between approximate and exact velocity profiles): for aspect ratio 1, the value of $a$ is 1.196; for aspect ratio 3, it is 1.188; and for aspect ratio 5, it is 1.184. Hence, the value $a=6/5=1.2$ we obtained by minimizing the residual, closely approximates the optimal error-minimizing values.

### v2

@@ Abstract
Over 125 years ago, Henry Selby Hele-Shaw realized that the depth-averaged flow in thin gap geometries can be closely approximated by two-dimensional (2D) potential flow, in a surprising marriage between the theories of viscous-dominated and inviscid flows. Hele-Shaw approximation allows visualization of potential flows over 2D airfoils and also undergirds important discoveries in the dynamics of interfacial instabilities and convection, yet it has found little use in modeling flows in microfluidic devices, although these devices often have thin gap geometries. Here, we derive a Hele-Shaw approximation for the flow in the kinds of thin gap geometries created within microfluidic devices. Using the Method of Weighted Residuals (MWR), we reinterpret the Hele–Shaw approximation as the leading term of an orthogonal polynomial expansion that can be systematically extended to higher-order corrections. The resulting leading-order equation coincides with the previously derived 2D approximations, but our derivation is shorter and more direct. By extending the expansion beyond leading order, we obtain a new reduced model that captures non-parabolic gap-wise velocity profiles and out-of-plane flow effects. We provide substantial numerical evidence showing that approximate equations can successfully model real microfluidic and inertial-microfluidic device geometries. By reducing three-dimensional (3D) flows to 2D models, our validated model will allow for accelerated device modeling and design.
@@ 1 Introduction
Although Hele-Shaw cells are often thought of as devices for engineering reduced complexity (2D) flows, thin gap domains are also commonly realized in microfluidic devices Stone et al. (2004); Di Carlo (2009), due to the nature of the fabrication methods used to create them. However, the Hele-Shaw approximation can not model no-slip boundary conditions on in-plane boundaries; it has found little use for dissecting the function of existing devices or designing new ones. It remains unanswered whether a different 2D model from Hele-Shaw exists, allowing the key dynamics present within a microfluidic or inertial microfluidic device to be predicted or understood, at far less computational cost than a fully 3D flow model.
The classical derivation of Hele-Shaw’s approximation assumes that, within the thin-gap geometry, the flow is planar but varies with $z$, the gap coordinate. Specifically assuming gap boundaries at $z=\pm\frac{h}{2}$, our putative velocity field is parabolic in $z$: $\mathbf{u}(x,y,z)\approx(u_{0}(x,y),v_{0}(x,y),0)\,6\,(\frac{1}{4}-\frac{z^{2}}{h^{2}})$, with $u_{0}$, $v_{0}$ the depth averaged velocities. Approximating: $\mu\nabla^{2}\mathbf{u}\approx\mu\partial^{2}\mathbf{u}/\partial z^{2}=-\frac{12\mu}{h^{2}}(u_{0}(x,y),v_{0}(x,y),0)$, and balancing with the pressure gradient, we arrive at a form of Darcy’s equation: $(u_{0},v_{0})=-\frac{h^{2}}{12\mu}\nabla p$, which models the pressure-driven flow of fluid through a porous medium, when velocities are averaged over multiple pores Whitaker (1986). Assuming that the flow is incompressible, $\nabla\cdot\mathbf{u}=0$, we obtain: $\nabla^{2}p=0$, so that the thin gap flow, $\mathbf{u}$, is a scalar potential flow, and $\frac{h^{2}p}{12\mu}$ the harmonic potential. Thus, despite the fact that flow between the plates is viscously dominated, the mean flows are captured by the same potential theory that models high-speed inviscid flows, such as those of ideal fluids over airfoils (Fig. 1(a,b)).
@@
Previous work has aimed to incorporate these additional effects into Hele-Shaw-type models. One approach involves adding boundary layer corrections Balsa (1998), though these calculations are often complex and highly geometry-dependent. Another direction is to modify the governing equations themselves. The Darcy-Brinkman equations add in-plane viscous stresses to the Hele-Shaw approximation, Brinkman (1949), and can be used for porous media Tsay & Weinbaum (1991); Howells (1974), as well as to thin-gap geometries with empirically determined relative coefficients on the in-plane and in-depth stresses Pepper et al. (2010). There are, by contrast, multiple pathways for incorporating inertial effects. When modeling porous media flow, a nonlinear pressure-flow relation (Darcy-Forchheimer) is commonly imposed Chen et al. (2001); Guta & Sundar (2010). Conversely, when Darcy’s law is arrived at by integrating the flow equations over some thin dimension, inertia-containing terms may likewise be integrated, though there are disagreements between papers on the precise coefficients thereby obtained Desvillettes et al. (2008); Inamdar et al. (2020).
@@
Here, we arrive at a Hele-Shaw-like approximation via a systematic expansion of the flow field in the gap coordinate. Our approximation naturally includes inertial and in-plane viscous stresses. Comparisons with full 3D simulations show that our approximation accurately models flows in geometries used in real microfluidic devices, including both conventional ($\mathrm{Re}=0$) and inertial ($\mathrm{Re}=1-100$) microfluidic flows.
2 A new perspective on the Hele-Shaw approximation
@@
We consider the Navier–Stokes equations for an incompressible, homogeneous fluid confined within a uniform-thickness gap. The equations are non-dimensionalized using the gap height $h$, so that the vertical coordinate $z$ lies in the interval 1/2<z<1/2-1/2<z<1/2. In the Hele–Shaw limit, the in-plane velocity components have a parabolic profile across the gap. The flow can be approximated by $\mathbf{u}(x,y,z)\approx 3/2\left(1-4z^{2}\right)\mathbf{u}_{0}(x,y)$, where $\mathbf{u}_{0}$ satisfies the no-slip boundary condition on $\mathbf{u}_{i}$ on any in-plane boundaries. Since $\int\limits_{-\frac{1}{2}}^{\frac{1}{2}}3/2\left(1-4z^{2}\right)\mathrm{d}z=1$, $u_{0}(x,y)$ is the depth-averaged velocity. Since all velocity components must vanish at the boundaries $z=\pm 1/2$, a natural choice for the higher order vertical basis functions is the family of Gegenbauer (ultraspherical) polynomials with parameter $\alpha=-1/2$, see Belinsky (2000). The Gegenbauer polynomials $C_{n}^{(\alpha)}$ are orthogonal with respect to the weight function $(1-z^{2})^{\alpha-1/2}$. In particular, $C_{n}^{(1/2)}$ reduces to the Legendre polynomial of degree $n$, and for $n\geq 2$ one has $C_{n}^{(-\frac{1}{2})}(z)=-\int\limits_{-1}^{z}C_{n-1}^{(\frac{1}{2})}(z)\mathrm{d}z$. The first few polynomials are
$C_{2}^{(-\frac{1}{2})}(z)=\frac{1}{2}\left(1-z^{2}\right),\;C_{3}^{(-\frac{1}{2})}(z)=\frac{1}{2}z\left(1-z^{2}\right),\;C_{4}^{(-\frac{1}{2})}(z)=\frac{1}{8}\left(5z^{2}-1\right)\left(1-z^{2}\right).$ (1)
Since $\lVert C_{n}^{(-\tfrac{1}{2})}(z)\rVert_{\infty}\sim n^{-\tfrac{3}{2}}$, we choose the basis as $a_{n}C_{n+2}^{(-\frac{1}{2})}(2z)$, where $a_{n}$ is a normalization factor chosen so that the basis remains of order one in magnitude. In addition, by the identity $(1-z^{2})C_{n}^{(\frac{3}{2})}(z)=(n+1)(n+2)C_{n+2}^{(-\frac{1}{2})}(z)$, it is therefore equivalent to choose $a_{n}(1-4z^{2})C_{n}^{\frac{3}{2}}(2z)$ as the basis, since this differs from our earlier choice only by a multiplicative constant.
Having established an orthogonal family of basis functions that satisfy the no-slip condition at $z=\pm 1/2$, we now represent the three-dimensional velocity field in this basis. Specifically, we expand
$\mathbf{u}=\sum_{n=0}^{\infty}\mathbf{u}_{n}(x,y)a_{n}C_{n+2}^{(-\frac{1}{2})}(2z),$ (2)
where $\mathbf{u}_{n}$ satisfies the no-slip boundary condition on any in-plane boundaries. The subsequent task in constructing the approximation is to specify the corresponding in-plane functions $\mathbf{u}_{n}(x,y)$ and derive the equations they must satisfy. As a first step, we illustrate the procedure by applying the series expansion to a unidirectional shear flow (rectilinear flow).
2.1 Poiseuille flow
@@
We first consider pressure-driven flow in a channel, aligned with the $x$-axis, with uniform rectangular cross-section $(y,z)\in[-L/2,L/2]\times[-1/2,1/2]$. Supposing an imposed pressure gradient of $G$, and scaling $p_{x}$ by $G$, and velocities by $Gh^{2}/(12\mu)$, the steady-state solution of Navier-Stokes equations takes the form $\mathbf{u}=(u(y,z),0,0)$, where $u$ satisfies a 2D Poisson equation $(\partial_{y}^{2}+\partial_{z}^{2})u=-12$. In the rectangular channel, this equation admits an exact solution (Cornish (1928); Boussinesq (1868)). Its series representation is:
$u_{exact}=\frac{3}{2}\left(1-4z^{2}\right)-\frac{48}{\pi^{3}}\sum\limits_{m=0}^{\infty}\frac{(-1)^{m}}{(1+2m)^{3}}\cos\left((2m+1)\pi z\right)\frac{\cosh\left((2m+1)\pi y\right)}{\cosh\left((2m+1)\pi\frac{L}{2}\right)}.$ (3)
By substituting the leading order forms of the components of $u\approx\frac{3}{2}(1-4z^{2})u_{0}(y)$ into the Poisson equation, we obtain the residual:
@@
where the first two terms come from $\partial_{y}^{2}+\partial_{z}^{2}$, and the third term is the dimensionless pressure gradient. Including additional bases such as $C_{3}^{(-\frac{1}{2})}(z),C_{4}^{(-\frac{1}{2})}(z),\ldots$ in our expansion would allow us to eliminate the $z$-dependent terms in Eq. (4), but at the expense of introducing a coupled system of equations for the unknown functions $u_{i}$. Before considering the more complex coupled system, we first seek an optimal closure involving only $u_{0}$, that is, a single equation for $u_{0}$ with constant coefficients. Therefore, we assume the optimal closure is of the form: $a\partial_{y}^{2}u_{0}(y)-12u_{0}(y)+12=0$ for some constant $a$ to be determined.  [^1] Alternatively, one may assume the more general form $a\partial_{y}^{2}u_{0}(y)-bu_{0}(y)+c=0$, but, by applying the residual minimization procedure introduced later, one would find that $b=12$ and $c=12$, which reduces exactly to the same equation as above. We can easily solve the equation and obtain the expression of $u_{0}(y)$:
$u_{0}(y)=1-\text{sech}\left(\frac{\sqrt{3}L}{\sqrt{a}}\right)\cosh\left(\frac{2\sqrt{3}y}{\sqrt{a}}\right).$ (5)
To determine $a$, we substitute it into Eq. (4), obtaining
$r(y,z;a)=u_{0}^{\prime\prime}(y;a)\left(\frac{3}{2}\left(1-4z^{2}\right)-a\right).$ (6)
We choose $a$ to minimize the second factor, integrating the squared residual against a weight function $w(z)$:
$E(a;w)\equiv\int\limits_{-\frac{1}{2}}^{\frac{1}{2}}\left(\frac{3}{2}\left(1-4z^{2}\right)-a\right)^{2}w(z)\mathrm{d}z.$ (7)
@@
$u\approx\frac{3}{2}(1-4z^{2})u_{0}(y)=\frac{3}{2}\left(1-4z^{2}\right)\left(1-\cosh\left(2\sqrt{3}y\right)\text{sech}\left(\sqrt{3}L\right)\right).$ (8)
@@
The key step to producing a better approximation is to observe that the space of functions spanned by Gegenbauer polynomials remains the natural space in which to optimize our approximation of Eq. (4). Since terms in Eq. (4) do not vanish at $z=\pm 1/2$, we multiply the equation by a weight $w(z)=(1-4z^{2})$, ensuring the product is contained within that space so that $r(z;a)w(z)=\sum_{n=0}^{\infty}\beta_{n}C_{n}^{(-\frac{1}{2})}(2z)$, for some set of coefficients $\beta_{m}$. Applying Parseval’s theorem to this series expansion gives:
$E(a;w)\equiv\int_{-1/2}^{1/2}r(z;a)^{2}(1-4z^{2})\mathrm{d}z=\int_{-1/2}^{1/2}(1-4z^{2})^{-1}\left(r(z;a)(1-4z^{2})\right)^{2}\mathrm{d}z$ (9)
$=\sum_{n=0}^{\infty}\beta_{n}^{2}\int\limits_{-\frac{1}{2}}^{\frac{1}{2}}(1-4z^{2})^{-1}C_{n}^{(-\frac{1}{2})}(2z)^{2}\mathrm{d}z.$
@@
Therefore, our optimal closure consists of minimizing Eq. (7) with $w(z)=(1-4z^{2})$. Under this weight, we obtain $E(a;w)=2a^{2}/3-8a/5+36/35$, which is minimized when $a=6/5$. This optimal closure leads to a different optimal profile:
@@
$u\approx\frac{3}{2}(1-4z^{2})u_{0}(y)=\frac{3}{2}\left(1-4z^{2}\right)\left(1-\cosh(\sqrt{10}y)\operatorname{sech}\left(L\sqrt{\frac{5}{2}}\right)\right).$ (10)
We provide two physical justifications for the parabolic weighting. First, the velocity field vanishes at the boundaries and attains its maximum at the center, so it is natural to assign greater weight near the center and less near the walls. Second, if the solution is independent of $y$, we have
$\partial_{z}^{2}u=f,\qquad u|_{z=\pm\tfrac{1}{2}}=0,$ (11)
with a general forcing term $f(z)$. The induced volume flux is then
$F=\int_{-\tfrac{1}{2}}^{\tfrac{1}{2}}u(y)\mathrm{d}z=\int_{-\tfrac{1}{2}}^{\tfrac{1}{2}}\tfrac{1}{8}\left(1-4z^{2}\right)f(z)\mathrm{d}z,$ (12)
where the second equality follows from integration by parts. Thus, when the error between the exact solution and its approximation is measured in terms of the resulting volume flux, the weight $1-4z^{2}$ naturally arises.
@@
Figure 2: Validating a 2D approximation for pressure-driven flow in a rectangular channel. (a) Velocity profile at $z=0$ for $L=1$. Red: Eq. (8); blue dashed: Eq. (10); orange dashed: higher order approximation from Eq. (26); black: exact solution. The inset shows a zoomed view of the interval $z\in[0,0.1]$. (b) Relative error in the $L^{\infty}$ norm, $\lVert u-u_{\text{app}}\rVert_{\infty}/\lVert u\rVert_{\infty}$, as a function of aspect ratio $L$. Red: Eq. (8); blue dashed: Eq. (10); orange dashed: Eq. (26). (c) Relative error in the $L^{2}$ norm, $\lVert u-u_{\text{app}}\rVert_{2}/\lVert u\rVert_{2}$, as a function of $L$. Red: Eq. (8); blue dashed: Eq. (10); orange dashed: Eq. (26); black dotted: $cL^{1/2}$, illustrating asymptotic scaling. (d) Relative error in the volume flux as a function of $L$. Red: Eq. (8); blue dashed: Eq. (10); orange dashed: Eq. (26).
Both approximations derived above are compared against the exact solution. The choice $a=6/5$ consistently yields a more accurate approximation across the entire $y$-domain (Fig. 2a) and for all tested values of $L$ (Fig. 2b). For example, when $L=1$, the relative errors for $a=1$ and $a=6/5$ are 0.115 and 0.031, respectively. The accuracy of the $a=6/5$ approximation is somewhat surprisingly, given that $L=1$ does not satisfy the classical scale-separation assumption $L\gg 1$. As $L$ increases beyond 1, the relative error decreases for both approximations. Though the error measured in the $L^{\infty}$ norm does not vanish in the infinite aspect ratio limit. This is because there always exists a narrow region near the side walls $y=\pm L/2$ where the vertical velocity profile deviates from a parabola due to the no-slip boundary condition. However, as the aspect ratio grows, this region occupies a smaller fraction of the domain, outside of which the pressure-driven flow in the channel approaches that in a two-parallel-plate geometry. When the error is measured in the $L^{2}$ norm instead, both approximations exhibit an error that decays as $L^{-1/2}$ in the infinite–aspect–ratio limit, as shown in Fig. 2(c).
Another quantity of interest is the volume flux of the approximate flow solution:
$F=\int\limits_{-1}^{1}\int\limits_{-\frac{L}{2}}^{\frac{L}{2}}u_{0}(y)Q_{0}(z)\mathrm{d}y\mathrm{d}z=L-\frac{\sqrt{a}\tanh\left(\frac{\sqrt{3}L}{\sqrt{a}}\right)}{\sqrt{3}}=L-\frac{\sqrt{a}}{\sqrt{3}}+\mathcal{O}\left(e^{-2\sqrt{\frac{3}{a}}L}\right),$ (13)
which may be compared to the exact flux formula obtained from (3)
$F=L-\frac{192}{\pi^{5}}\sum\limits_{m=0}^{\infty}\frac{\tanh\left(\frac{L\pi}{2}(2m+1)\right)}{(2m+1)^{5}}=L-\frac{186\zeta(5)}{\pi^{5}}+\mathcal{O}\left(e^{-L\pi}\right),$ (14)
where $\zeta(s)$ is the Riemann zeta function. Figure 2(d) shows the relative error in the volume flux computed using Eq. (8) and Eq. (10), demonstrating that the latter achieves consistently better accuracy. For reference, if side-wall effects are ignored, the solution reduces to $u(y,z)=\tfrac{3}{2}(1-4z^{2})$, which yields a flux of $F=L$. In terms of the volume flux, Eq. (13) shows that a channel of length $L$ with no-slip side walls is effectively equivalent to a channel of length $L-\sqrt{2/5}\approx L-0.63246$ with slip side walls. This agrees closely with the asymptotic exact flux, $L-\tfrac{186\zeta(5)}{\pi^{5}}\approx L-0.63025$.
To probe the optimality of our closure relation more closely, at each aspect ratio, we numerically compute the value of $a$ that minimizes
$\frac{\lVert u_{exact}(y,z)-\frac{3}{2}\left(1-4z^{2}\right)u_{0}(y;a)\rVert_{\infty}}{\lVert u_{exact}(y,z)\rVert_{\infty}}.$ (15)
For aspect ratio 1, the value of $a$ is 1.196; for aspect ratio 3, it is 1.188; and for aspect ratio 5, it is 1.184. Hence, the value $a=6/5=1.2$ we obtained by minimizing the residual, closely approximates the optimal error-minimizing values.
We remark that minimizing (15) requires knowledge of the exact solution to $\Delta u=-12$. For general or more complicated problems, this is impractical. If the exact solution were available, constructing such an approximation would be of limited interest. By contrast, the $L^{2}$ minimization in (7) does not require solving the governing equation in full, nor even computing $u_{0}(y;a)$. Interestingly, although $u_{0}(y;a)$ depends on $a$, the value of $a$ obtained by minimizing $E(a;w)$ does not necessarily minimize the entire residual $r(y,z;a)$ in the $L^{\infty}$ norm or under other metrics. Nevertheless, it provides a good approximation to the solution. Because this approach yields a reasonably accurate result without requiring the solution of any differential equation, it is computationally attractive.
2.2 Method of weighted residuals
Our closure method follows the general framework of the method of weighted residuals (MWR) Finlayson (2013), which has been applied to derive reduced-order approximation for thin gap fluid flows ranging from free-surface thin-film flows Kalliadasis et al. (2011); Ruyer-Quil & Manneville (2000), Hele–Shaw flows Ruyer-Quil (2001), to the thermal convection in a vertical slot Papanicolaou et al. (2009).